# Covariate-Adaptive Randomization in Clinical Trials without Inflated Variances

### v2

@@ 4.1 The properties of the CAR procedures
$=$ $-r\left|\ell_{n+1}-\rho\right|\cdot\left|\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\right|\cdot\|\bm{\Lambda}_{n}\|^{r-2}+c_{r}\left(\|\bm{\phi}_{n+1}\|^{r}+\|\bm{\phi}_{n+1}\|^{2}\|\bm{\Lambda}_{n}\|^{r-2}\right).$
Notice that there is a $\delta>0$, $|\ell(x)-\rho|\geq|x||l^{\prime}(0)|/2$ when $|x|\leq\delta$. While, when $|x|\geq\delta$, $|\ell(x)-\rho|\geq(\rho-\ell(\delta)\wedge(\ell(-\delta)-\rho)>0$. It follows that, there exists a contact $\lambda_{1}>0$ such that $|\ell(x)-\rho|\geq\lambda_{1}(|x|\wedge 1)$. Thus, $\textsf{E}\left[\left|\ell_{n+1}-\rho\right|\cdot\left|\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\right|\big|\mathscr{F}_{n}\right]\geq\lambda_{1}n^{\gamma}\hbar(\bm{\lambda}_{n}/n^{\gamma})$, where $\hbar(\bm{x})=\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]$. Next, we show that here exists a constant $\lambda_{2}>0$ such that $\hbar(\bm{x})\geq\lambda_{2}(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)$. Under the assumption that $\Sigma=\textsf{E}[\bm{\phi}^{\prime}\bm{\phi}]$ is non-singular, we have $\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]\neq 0$ for all $\bm{x}\neq\bm{0}$. In fact, if $\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]=0$, then $\textsf{P}(\langle\bm{x},\bm{\phi}\rangle=0)=1$ which implies that $\bm{x}\Sigma\bm{x}^{\prime}=\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\right]=0$. Thus, we must have $\bm{x}=\bm{0}$ since $\Sigma$ is non-singular.
Notice that $\hbar(\bm{x})$ is a continuous function. Let $\lambda_{0}=\inf\{\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]:\|\bm{x}\|=1\}$. Then $\lambda_{0}>0$. Since $\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|\leq(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)(\|\bm{\phi}\|^{2}+\|\bm{\phi}\|)$, by the dominated convergence theorem, as $\|\bm{x}\|\to 0$ with $\bm{x}/\|\bm{x}\|\to\bm{x}_{0}$,
$\lim_{\|\bm{x}\|\to 0}\frac{\hbar(\bm{x})}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}=$ $\textsf{E}\left[\lim_{\|\bm{x}\|\to 0}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}\right]=\textsf{E}\left[\lim_{\|\bm{x}\|\to 0}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}}{\|\bm{x}\|^{2}}\right]$
$=$ $\textsf{E}\left[\left|\langle\bm{x}_{0},\bm{\phi}\rangle\right|^{2}\right]\geq\big(\textsf{E}\left[\left|\langle\bm{x}_{0},\bm{\phi}\rangle\right|\right]\big)^{2}=\lambda_{0}^{2}>0;$
and, as $\|\bm{x}\|\to\infty$ with $\bm{x}/\|\bm{x}\|\to\bm{x}_{0}$,
$\lim_{\|\bm{x}\|\to\infty}\frac{\hbar(\bm{x})}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}=$ $\textsf{E}\left[\lim_{\|\bm{x}\|\to\infty}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}\right]=\textsf{E}\left[\lim_{\|\bm{x}\|\to\infty}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|}{\|\bm{x}\|}\right]$
$=$ $\textsf{E}\left[\left|\langle\bm{x}_{0},\bm{\phi}\rangle\right|\right]\geq\lambda_{0}>0.$
Thus, there exists $\lambda_{2}>0$ such that $\hbar(\bm{x})\geq\lambda_{2}(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)$. It follows that
@@
where $\beta_{r}=\textsf{E}[\|\bm{\phi}(\bm{X})\|^{r}]$. Hence,
@@
$\leq$ $-r\lambda_{1}\lambda_{2}\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)\right]+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-2}\right]\right)$ (4.32)
$\leq$ $c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-2}\right]\right).$ (4.33)
Taking $r=2$ in (4.33) yields $\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]\leq 2c_{2}\beta_{2}n$. Thus, $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq(\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}])^{r/2}\leq C_{r}n^{r/2}$ for all $n$ and $r\in(0,2]$. By (4.33) and the induction,
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq C_{r}n^{r/2},\;\;n\geq 1,r\in(0,r_{0}].$ (4.34)
Let $E=\{\|\bm{\Lambda}_{n}\|\leq n^{\gamma}\}$ and $Q_{n,r}=\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)\right]$. Then
@@
For $\frac{1}{p}+\frac{1}{q}=1$, $p,q>1$, by the Hölder inequality we have
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}I_{E^{c}}\right]=\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{\frac{r-1}{p}}\|\bm{\Lambda}_{n}\|^{\frac{r-1}{q}+1}I_{E^{c}}\right]$
$\leq$ $\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}I_{E^{c}}\right]\right)^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+q}I_{E^{c}}\right]\right)^{1/q}\leq Q_{n,r}^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+q}\right]\right)^{1/q}.$
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq n^{\gamma}Q_{n,r}+Q_{n,r}^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+q}\right]\right)^{1/q}.$
@@
$r\lambda_{1}\lambda_{2}Q_{m,r}+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)\geq 0.$
@@
$Q_{m,r}\leq C_{r}\left(1+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)$
@@
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m+1}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r}\right]+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)$
$\leq$ $m^{\gamma}Q_{m,r}+Q_{m,r}^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1+q}\right]\right)^{1/q}+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)$
$\leq$ $C_{r}\left(m^{\gamma}\left(1+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)+\left(1+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)^{1/p}\left(\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1+q}]\right)^{1/q}\right).$ (4.35)
Proof of Theorem 2.1. Let $r=2$ and $r-1+q=r_{0}$. Notice $\textsf{E}[\|\bm{\Lambda}_{m}\|^{r_{0}}]\leq C_{r_{0}}m^{r_{0}/2}$. We obtain
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]\leq C\left(n^{\gamma}+n^{1+\frac{2-r_{0}}{2(r_{0}-1)}}\right)=o(n),$
by (4.1). Thus, (2.5) is proven.
Notice that
$\textsf{E}|(\ell_{n}-1)Z_{n}|\leq$ $cM\textsf{E}\left[\frac{\|\bm{\Lambda}_{n-1}\|\cdot\|\bm{\phi}_{n}\|}{(n-1)^{\gamma}}\right]+\textsf{E}[|Z|I\{|Z|>M\}]$
@@
$\to$ $0\text{ as }n\to\infty\text{ and then }M\to\infty,$
since
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]=O\left(n^{\gamma/2}+n^{\frac{1}{4}+\frac{1}{4(r_{0}-1)}}\right)=o(n^{\gamma}),\;\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}.$
(2.6) is proven.
For (2.7), let $\Delta M_{i,Z}=:(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]$, $M_{n,Z}=\sum_{i=1}^{n}\Delta M_{i,Z}$. Notice that $\{\Delta M_{i,Z},\mathscr{F}_{i},i=1,2,\ldots\}$ is a sequence of martingale differences. Since $|\Delta M_{i,Z}|\leq|Z_{i}|+\textsf{E}|Z|$, $\{\Delta M_{i,Z},i=1,2,\ldots\}$ is uniformly integrable. Thus, $\frac{1}{n}\textsf{E}[|M_{n,Z}|\to 0$ (Hall and Heyde, , 1980, Theorem 2.2). On the other hand,
@@
For (2.8), we denote $\Delta M_{i,\widetilde{Z}}=:(T_{i}-\rho)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]$. Suppose $r_{0}\geq 4$ and $\textsf{E}[Z^{2}]<\infty$. Then
@@
$=$ $\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]-\ell^{\prime}(0)h_{\widetilde{Z}}\Big(\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\Big)\right|\leq C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{2}\|\bm{\Phi}\|^{2}|\widetilde{Z}|]}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]$
@@
Also, since $\bm{x}_{0}\Sigma\bm{x}_{0}=\textsf{E}(\textsf{Pj}[Z|\bm{\phi}])^{2}\leq\textsf{E}Z^{2}$, we have $\|\bm{x}_{0}\|\leq C_{0}(\textsf{E}Z^{2})^{1/2}$. Hence, we have
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+\theta_{n}\left(\|\bm{\Lambda}_{n}\|+\sum_{i=2}^{n-1}\frac{\|\bm{\Lambda}_{i}\|^{2}}{i^{2\gamma}}\right),\;\;|\theta_{n}|\leq C_{0},$ (4.38)

### v3

@@ 4.1 The properties of the CAR procedures
$=$ $-r\left|\ell_{n+1}-\rho\right|\cdot\left|\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\right|\cdot\|\bm{\Lambda}_{n}\|^{r-2}+c_{r}\left(\|\bm{\phi}_{n+1}\|^{r}+\|\bm{\phi}_{n+1}\|^{2}\|\bm{\Lambda}_{n}\|^{r-2}\right),$
where the last equality is due to the fact that $\ell_{n+1}$ is a non-increasing function of $\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle$. Notice that there is a $\delta>0$, $|\ell(x)-\rho|\geq|x||\ell^{\prime}(0)|/2$ when $|x|\leq\delta$. While, when $|x|\geq\delta$, $|\ell(x)-\rho|\geq(\rho-\ell(\delta))\wedge(\ell(-\delta)-\rho)>0$. It follows that, there exists a constant $\lambda_{1}>0$ such that $|\ell(x)-\rho|\geq\lambda_{1}(|x|\wedge 1)$. Thus, $\textsf{E}\left[\left|\ell_{n+1}-\rho\right|\cdot\left|\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\right|\big|\mathscr{F}_{n}\right]\geq\lambda_{1}n^{\gamma}\hbar(\bm{\Lambda}_{n}/n^{\gamma})$, where $\hbar(\bm{x})=\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]$.
Let $\lambda_{min}$ be the smallest eigenvalue of $\Sigma$. Then $\textsf{E}\langle\bm{x},\bm{\phi}\rangle^{2}=\bm{x}\Sigma\bm{x}^{\prime}\geq\lambda_{min}\|\bm{x}\|^{2}$. Choose $M\geq 1$ large enough such that $\textsf{E}[\|\bm{\phi}\|^{2}I\{\|\bm{\phi}\|\geq M\}]\leq\lambda_{min}/2$. Notice
$\langle\bm{x},\bm{\phi}\rangle^{2}\leq$ $\langle\bm{x},\bm{\phi}\rangle^{2}I\{|\langle\bm{x},\bm{\phi}\rangle|\leq 1\}$
$+M\|\bm{x}\|\,\big|\langle\bm{x},\bm{\phi}\rangle\big|I\{|\langle\bm{x},\bm{\phi}\rangle|>1\}+\|\bm{x}\|^{2}\|\bm{\phi}\|^{2}I\{|\|\bm{\phi}\||\geq M\}$
$\leq$ $[1\vee(M\|\bm{x}\|)]\left[\langle\bm{x},\bm{\phi}\rangle^{2}\wedge|\langle\bm{x},\bm{\phi}\rangle|\right]+\|\bm{x}\|^{2}\|\bm{\phi}\|^{2}I\{|\|\bm{\phi}\||\geq M\}.$
It follows that
$\hbar(\bm{x})\geq\frac{\textsf{E}\langle\bm{x},\bm{\phi}\rangle^{2}-\|\bm{x}\|^{2}\textsf{E}[\|\bm{\phi}\|^{2}I\{\|\bm{\phi}\|\geq M\}]}{1\vee(M\|\bm{x}\|)}\geq\frac{\lambda_{min}\|\bm{x}\|^{2}/2}{1\vee(M\|\bm{x}\|)}.$
Let $\lambda_{2}=\lambda_{min}/(2M)$. We have $\hbar(\bm{x})\geq\lambda_{2}(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)$. It follows that
@@
where $\beta_{r}=\textsf{E}[\|\bm{\phi}(\bm{X})\|^{r}]$. Consider the function
$f(x)=-\alpha_{1}x^{r-1}\left(\frac{x}{n^{\gamma}}\wedge 1\right)+\alpha_{2}x^{r-2},\;x\geq 0,$
where $\alpha_{1},\alpha_{2}>0$. When $0\leq x\leq n^{\gamma}$,
$f(x)=-\alpha_{1}\frac{x^{r}}{n^{\gamma}}+\alpha_{2}x^{r-2}\leq$ $\begin{cases}0,&\text{ when }\alpha_{1}\frac{x^{2}}{n^{\gamma}}\geq\alpha_{2}\\ \alpha_{2}\Big(\frac{\alpha_{2}}{\alpha_{1}}\Big)^{\frac{r-2}{2}}n^{\frac{\gamma}{2}(r-2)},&\text{ when }\alpha_{1}\frac{x^{2}}{n^{\gamma}}<\alpha_{2}\end{cases}$
$\leq$ $\alpha_{2}\Big(\frac{\alpha_{2}}{\alpha_{1}}\Big)^{\frac{r-2}{2}}n^{\frac{\gamma}{2}(r-2)}.$
Similarly, when $x>n^{\gamma}$,
$f(x)=-\alpha_{1}x^{r-1}+\alpha_{2}x^{r-2}\leq\alpha_{2}\Big(\frac{\alpha_{2}}{\alpha_{1}}\Big)^{r-2},\;x\geq 0.$
It follows that
$f(x)\leq\alpha_{2}\Big(\frac{\alpha_{2}}{\alpha_{1}}\Big)^{r-2}+\alpha_{2}\Big(\frac{\alpha_{2}}{\alpha_{1}}\Big)^{\frac{r-2}{2}}n^{\frac{\gamma}{2}(r-2)}.$
By choosing $\alpha_{1}=r\lambda_{1}\lambda_{2}/2$ and $\alpha_{2}=c_{r}\beta_{2}$, it follows that
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}\Big(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\Big)+C_{r}n^{\frac{\gamma}{2}(r-2)}.$
Hence, we have the following recursive formula:
@@
$\leq$ $-\frac{r\lambda_{1}\lambda_{2}}{2}\textsf{E}\Big[\|\bm{\Lambda}_{n}\|^{r-1}\Big(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\Big)\Big]+C_{r}n^{\frac{\gamma}{2}(r-2)}$ (4.34)
$\leq$ $C_{r}n^{\frac{\gamma}{2}(r-2)}.$ (4.35)
By (4.35),
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r_{0}}]\leq C_{r_{0}}n^{\frac{\gamma}{2}(r_{0}-2)+1},\;\;n\geq 1.$ (4.36)
Let $E=\{\|\bm{\Lambda}_{n}\|\leq n^{\gamma}\}$ and $Q_{n,r}=\textsf{E}\big[\|\bm{\Lambda}_{n}\|^{r-1}\big(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\big)\big]$. Then
@@
For $\frac{1}{f}+\frac{1}{g}=1$, $f,g>1$, by the Hölder inequality we have
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}I_{E^{c}}\right]=\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{\frac{r-1}{f}}\|\bm{\Lambda}_{n}\|^{\frac{r-1}{g}+1}I_{E^{c}}\right]$
$\leq$ $\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}I_{E^{c}}\right]\right)^{1/f}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+g}I_{E^{c}}\right]\right)^{1/g}\leq Q_{n,r}^{1/f}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+g}\right]\right)^{1/g}.$
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq n^{\gamma}Q_{n,r}+Q_{n,r}^{1/f}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+g}\right]\right)^{1/g}.$
@@
$\frac{r\lambda_{1}\lambda_{2}}{2}Q_{m,r}+C_{r}m^{\frac{\gamma}{2}(r-2)}\geq 0.$
@@
$Q_{m,r}\leq\frac{2C_{r}}{r\lambda_{1}\lambda_{2}}m^{\frac{\gamma}{2}(r-2)},$
@@
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m+1}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r}\right]+C_{r}m^{\frac{\gamma}{2}(r-2)}$
$\leq$ $m^{\gamma}Q_{m,r}+Q_{m,r}^{1/f}\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1+g}\right]\right)^{1/g}+C_{r}m^{\frac{\gamma}{2}(r-2)}$
$\leq$ $C_{r}\left[m^{\frac{\gamma}{2}r}+\big(m^{\frac{\gamma}{2}(r-2)}\big)^{1/f}\left(\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1+g}]\right)^{1/g}\right].$
It follows that
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq C_{r}\Big[n^{\frac{\gamma}{2}r}+\big(n^{\frac{\gamma}{2}(r-2)}\big)^{1/f}\Big(\max_{m\leq n}\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1+g}]\Big)^{1/g}\Big],\;r\in[2,r_{0}].$ (4.37)
When $r\in[0,2]$, we have the elementary inequalities that
$\|\bm{u}+\bm{v}\|^{r}-\|\bm{u}\|^{r}\leq r\langle\bm{u},\bm{v}\rangle\|\bm{u}\|^{r-2}+\frac{r}{2}\|\bm{v}\|^{2}\|\bm{u}\|^{r-2},$
and $\|\bm{u}+\bm{v}\|^{r}-\|\bm{u}\|^{r}\leq 2\|\bm{v}\|^{r}+\|\bm{u}\|^{r}.$ We use the first inequality when $\|\bm{u}\|\geq 1$ and the second inequality when $\|\bm{u}\|\leq 1$. It follows that
$\|\bm{u}+\bm{v}\|^{r}-\|\bm{u}\|^{r}\leq$ $r\langle\bm{u},\bm{v}\rangle\|\bm{u}\|^{r-2}I\{\|\bm{u}\|\geq 1\}+\frac{r}{2}\|\bm{v}\|^{2}+2\|\bm{v}\|^{r}+1$
$\leq$ $r\langle\bm{u},\bm{v}\rangle\|\bm{u}\|^{r-2}+\|\bm{v}\|^{2}+2\|\bm{v}\|^{r}+2\|\bm{v}\|+1.$
Similar to (4.1), we have
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-r\lambda_{1}\lambda_{2}\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)+C_{r},\;r\in[1,2].$ (4.38)
Thus,
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\right]-\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq-r\lambda_{1}\lambda_{2}\textsf{E}\Big[\|\bm{\Lambda}_{n}\|^{r-1}\Big(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\Big)\Big]+C_{r},\;r\in[1,2].$
With this inequality taking the place of (4.34), we can show that
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq C_{r}\Big[n^{\gamma}+\Big(\max_{m\leq n}\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1+g}]\Big)^{1/g}\Big],\;r\in[1,2].$ (4.39)
Proof of Theorem 2.1. (i) Let $r-1+g=r_{0}$. Notice (4.36). We obtain
$\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\right)^{1/r}\leq$ $C_{r,r_{0}}\left(n^{\frac{\gamma}{2}r}+n^{\frac{\gamma(r-2)}{2f}+\frac{(r_{0}-2)\gamma+2}{2g}}\right)^{1/r}$
$\leq$ $C_{r,r_{0}}\left(n^{\frac{\gamma}{2}}+n^{\frac{\gamma}{2}-\frac{(r_{0}+2-r)\gamma-2}{2r(r_{0}+1-r)}}\right),\;r\in[2,r_{0}],$
by (4.37), and
$\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\right)^{1/r}\leq$ $C_{r,r_{0}}\left(n^{\gamma}+n^{\frac{(r_{0}-2)\gamma+2}{2g}}\right)^{1/r}\leq C_{r,r_{0}}\left(n^{\frac{\gamma}{r}}+n^{\frac{(r_{0}-2)\gamma+2}{2r(r_{0}+1-r)}}\right),\;r\in[1,2],$
by (4.39). Thus, (2.5) and (2.6) are proven. Taking $r=2$ and $r=r_{0}$ in (2.5) respectively yields (2.7).
(ii) For (2.8), if $\gamma>\frac{2}{3r_{0}-2}$, then
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]\leq\Big(\textsf{E}\|\bm{\Lambda}_{n}\|^{2}\Big)^{1/2}=O\left(n^{\gamma/2}+n^{\frac{(r_{0}-2)\gamma+2}{4(r_{0}-1)}}\right)=o(n^{\gamma}),$
by (2.7). If $\gamma>\frac{4}{r_{0}^{2}+5}$ and $2<r_{0}\leq 3$, then by taking $r=(r_{0}+1)/2$ in (2.6), we have
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]\leq\Big(\textsf{E}\|\bm{\Lambda}_{n}\|^{r}\Big)^{1/r}=O\Big(n^{\frac{\gamma}{r}}+n^{\gamma+\frac{4-\gamma(r_{0}^{2}+5)}{(r_{0}+1)^{2}}}\Big)=o(n^{\gamma}).$
If $\gamma>\frac{8}{(r_{0}+2)^{2}+4}$ and $r_{0}\geq 4$, then by taking $r=r_{0}/2$ in (2.5), we have
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]\leq\Big(\textsf{E}\|\bm{\Lambda}_{n}\|^{r}\Big)^{1/r}=O\Big(n^{\frac{\gamma}{2}}+n^{\gamma+\frac{2-\gamma((r_{0}+2)^{2}+4)/4}{r_{0}(r_{0}/2+1)}}\Big)=o(n^{\gamma}).$
Notice that $|\ell(x)-\rho|\leq c|x|$ and $|\ell(x)-\rho|\leq 1$. Thus,
$\textsf{E}|(\ell_{n}-\rho)Z_{n}|\leq$ $cM\textsf{E}\left[\frac{\|\bm{\Lambda}_{n-1}\|\cdot\|\bm{\phi}_{n}\|}{(n-1)^{\gamma}}\right]+\textsf{E}[|Z|I\{|Z|>M\}]$
@@
$\to$ $0\text{ as }n\to\infty\text{ and then }M\to\infty.$
For (2.9), let $\Delta M_{i,Z}=:(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]$, $M_{n,Z}=\sum_{i=1}^{n}\Delta M_{i,Z}$. Notice that $\{\Delta M_{i,Z},\mathscr{F}_{i},i=1,2,\ldots\}$ is a sequence of martingale differences. Since $|\Delta M_{i,Z}|\leq|Z_{i}|+\textsf{E}|Z|$, $\{\Delta M_{i,Z},i=1,2,\ldots\}$ is uniformly integrable. Thus, $\frac{1}{n}\textsf{E}|M_{n,Z}|\to 0$ (Hall and Heyde, , 1980, Theorem 2.2). On the other hand,
@@
(iii) For ((iii)), we denote $\Delta M_{i,\widetilde{Z}}=:(T_{i}-\rho)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]$. Suppose $r_{0}\geq 4$ and $\textsf{E}[Z^{2}]<\infty$. Then
@@
$=$ $\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]-\ell^{\prime}(0)h_{\widetilde{Z}}\Big(\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\Big)\right|\leq C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{2}\|\bm{\phi}\|^{2}|\widetilde{Z}|}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]$
@@
Also, since $\bm{x}_{0}\Sigma\bm{x}_{0}^{\prime}=\textsf{E}(\textsf{Pj}[Z|\bm{\phi}])^{2}\leq\textsf{E}Z^{2}$, we have $\|\bm{x}_{0}\|\leq\lambda_{min}^{-1/2}(\textsf{E}Z^{2})^{1/2}$. Hence, we have
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+\theta_{n}(\textsf{E}Z^{2})^{1/2}\left(\|\bm{\Lambda}_{n}\|+\sum_{i=2}^{n-1}\frac{\|\bm{\Lambda}_{i}\|^{2}}{i^{2\gamma}}\right),\;\;|\theta_{n}|\leq C_{0},$ (4.42)